OpenStudy (anonymous):

how to write ''passes through the points (-3,6),(4,-1) in standard form?

OpenStudy (anonymous):

f(-3)=6 and f(4)=-1

Directrix (directrix):

(-3,6),(4,-1) I assume the task is to write the equation of the line determined by these two points and then to write that equation in standard form for the equation of a line. In y = mx + b form, the equation of that line is y = -x + 3. Standard Form: the standard form of a line is in the form Ax + By = C where A is a positive integer, and B, and C are integers. So, y = -x + 3 becomes 1*x + (1)*y = 3 in standard form where A = 1, B = 1, and C = 3. @Lauryn29611